Business and finance. Jose has savings accounts for each of his five children. They contain $215, $156, $318, $75, and $25. Find the median and the mean amount of money per account.
Answer by praseenakos@yahoo.com(507)


Mean = sum of all the terms/ number of terms
=($25+$75+$156+$215+$318)/5
= (789)/5
= 157.8
Median= (n+1)/2 term
=(5+1)/2 term
= 6/2 term
= 3rd term
therefore Median = 156 (arrange the terms in the ascending order
and then choose the third term)